# STaRK: Benchmarking LLM Retrieval on Textual and Relational Knowledge Bases

### v2

@@ Abstract
Answering real-world complex queries, such as complex product search, often requires accurate retrieval from semi-structured knowledge bases that involve blend of unstructured (e.g., textual descriptions of products) and structured (e.g., entity relations of products) information. However, previous works have mostly studied textual and relational retrieval tasks as separate topics. To address the gap, we develop STaRK, a large-scale Semi-structure retrieval benchmark on Textual and Relational Knowledge Bases. Our benchmark covers three domains/datasets: product search, academic paper search, and queries in precision medicine. We design a novel pipeline to synthesize realistic user queries that integrate diverse relational information and complex textual properties, together with their ground-truth answers (items). We conduct rigorous human evaluation to validate the quality of our synthesized queries. We further enhance the benchmark with high-quality human-generated queries to provide an authentic reference. STaRK serves as a comprehensive testbed for evaluating the performance of retrieval systems driven by large language models (LLMs). Our experiments suggest that STaRK presents significant challenges to the current retrieval and LLM systems, indicating the demand for building more capable retrieval systems. The benchmark data and code are available on https://github.com/snap-stanford/stark.
@@
Table 1: Example queries from STaRK, which involve semi-structured information, where the relational and textual aspects are highlighted.
@@ 1 Introduction
To answer such queries the underlying knowledge can be represented in semi-structured knowledge bases (SKBs) [31, 43, 34], which integrate unstructured data, such as textual descriptions and natural language expressions (e.g., description of the tricycle), with structured data, like entity interactions on knowledge graph (e.g., a tricycle “brand” is Radio Flyer). This allows the SKBs to represent comprehensive knowledge in specific applications, making them indispensable in domains such as e-commerce [13], social media [28], and precision medicine [6, 16, 21].
@@
Limitations of prior works. Prior works and benchmarks have primarily focused on either purely textual queries on unstructured knowledge [18, 12, 23, 26, 46] or structured SQL or knowledge graph queries [2, 50, 48, 38], which are inadequate to study the complexities of retrieval tasks involving SKBs. Another line of works [10, 22, 48, 44, 45, 39] have explored the middle ground via multi-hop reasoning within single or across multiple documents. However, they are either limited in the span of knowledge or lack of diverse textual properties and explicit relational information. Recently, large language models (LLMs) have demonstrated significant potential on information retrieval tasks [54, 12, 27, 36]. Nevertheless, it remains an open question how effectively LLMs can be applied to the specific challenge of retrieval from SKBs. It is also important to note that existing works focus mainly on general knowledge, e.g., from Wikipedia. In fact, the knowledge may commonly come from private sources, requiring information retrieval systems to operate on private SKBs. Therefore, there is a gap in our understanding of how current LLM retrieval systems handle the complex interplay between textual and relational requirements in queries that optionally involve private knowledge.
@@
Present work. To address this gap, we present a large-scale Semi-structure retrieval benchmark on Textual and Relational Knowledge Bases (STaRK) (Figure 1). The key technical challenge that we solve is how to accurately simulate user queries on SKBs. This difficulty arises from the interdependence of textual and relational information, which can lead to challenges in precisely filtering through millions of candidates to construct the ground-truth answers. Additionally, ensuring that queries are relevant to real-world applications and resembles real-world scenarios adds further complexity to the benchmarking process.
@@
We develop a novel pipeline that simulates user queries and constructs precise ground truth answers using three SKBs built from extensive texts and millions of entity relations from public sources. We validate the quality of queries in our benchmark through detailed analysis and human evaluation, focusing on their naturalness, diversity, and practicality. Furthermore, we incorporate 274 human-generated queries into our benchmark to compare with synthesized queries and enrich the testing scenarios. With STaRK, we delve deeper into retrieval tasks on SKBs, evaluate the capability of current retrieval systems, and provide insights for future advancement. Key features of STaRK are:
@@
  • Natural-sounding queries on SKBs (Table 1): Queries in our benchmark are crafted to incorporate rich relational information and complex textual properties. Additionally, these queries closely mirror the types of questions users would naturally ask in real-life scenarios, e.g., with flexible query formats and possibly with extra contexts.
@@
  • Diverse domains: Our benchmark spans a wide range of domains, reflecting the diverse nature of real-world information needs. It covers three different knowledge bases1footnotetext: 1 We provide an interactive interface for exploring the SKBs at https://huggingface.co/spaces/snap-stanford/SKB-Explorer with applications on product recommendation, academic paper search, and precision medicine inquiry. Therefore, STaRK provides a comprehensive evaluation of retrieval systems across various contexts and knowledge domains.
We also perform extensive experiments on each benchmark dataset for the prevailing retrieval systems. We highlight research challenges especially on the capability to handle textual and relational requirements and the latency of retrieval systems on large-scale SKBs that involve million-scale entities or relations. Finally, we provide insights into future directions to build more capable retrieval systems over private knowledge bases.
@@ 2 Benchmarking Retrieval Tasks over Textual and Relational Knowledge
2.1 Problem Definition
We are given a Semi-Structured Knowledge Base (SKB), which consists of a knowledge graph $G$ and a collection of free text documents $D$. Formally, let $G=(V,E)$ be the knowledge graph, where $V$ is the set of nodes and $E\subseteq V\times V$ is the set of edges representing relationships between nodes. $D=\bigcup_{i\in V}D_{i}$ be the collection of free-form text documents associated with the nodes, where $D_{i}$ is the set of documents associated with node $i$. For example, the product knowledge graph in e-commerce can capture relationships between products and brands/colors/categories, and the corresponding text documents include product descriptions, reviews, etc..
2.2 Semi-structured Knowledge Bases (SKBs)
@@
As shown in Figure 2, we construct three SKBs from the relational structure between entities and the textual information associated with a subset of the entities. We present the statistics of the relational structure in Table 2 and introduce each SKB as follows:
@@
Amazon Semi-structured Knowledge Base. This knowledge base is derived from the Sports and Outdoors category from Amazon Product Reviews [13] and Amazon Question and Answer Data [29]. The knowledge base features four entity types: product, brand, color, and category, and four types of relational information: also_bought, also_viewed between product entities, has_brand, has_color, has_category relations between product and brand, color, and category entities, respectively. In total, it comprises around 1.0M entities (product entities: 0.9M, brand entities: 0.1M, category entities: 1.4k, color entities: 1.7k) and 9.4M relations (also_bought: 2.8M, also_viewed: 1.9M, has_brand: 1.7M, has_category: 2.3M, has_color: 0.6M). We obtain the textual information by combining the meta data from Amazon Product Reviews with the customer Q&A records from Amazon Question and Answer Data. This provides a rich amount of textual data, including product titles, descriptions, prices, customer reviews, and related customer Q&A for each product. For other entities, we extract their names or titles as the textual attribute. Amazon SKB features an extensive textual data, which is largely contributed from customer reviews and Q&A.
@@
MAG Semi-structured Knowledge Base. This knowledge base is constructed from obgn-MAG [17], obgn-papers100M [17], and Microsoft Academic Graph (version 2019-03-22) [43, 37]. The knowledge base contains around 1.9M entities under four entity types (author: 1.1M, paper: 0.7M, institution: 8.7K, field_of_study: 59.5k) and 39.8M relations under four relation types (author_writes_paper: 13.5M, paper_has_field_of_study: 14.5M, paper_cites_paper: 9.7M, author_affiliated_with_institution: 2.0M). We construct the knowledge base by selecting papers1 shared between obgn-MAG and obgn-papers100M. The relational structure comes from the extracted subgraph in obgn-mag based on the selected papers. The textual information such as titles and abstracts is sourced from obgn-papers100M. Additionally, we augment the textual data by integrating details from the Microsoft Academic Graph database, providing extra information like paper venue, author and institution names. This SKB demonstrates a large number of entities and relations associate with paper nodes, especially on citation and authorship relation types.
@@
Prime Semi-structured Knowledge Base. We leverage the existing knowledge graph PrimeKG [6] which contains ten entity types including disease, drug, gene/protein, and eighteen relation types, such as associated_with, synergistic_interaction, indication, expression_present. The entity count in our knowledge base is approximately 129.3K, with around 8.1M relations. The details on entity numbers and relation tuples are available in Appendix B.3. Compared to the Amazon and MAG SKBs, Prime SKB is denser and features a greater variety of relation types. While PrimeKG already provides text information on disease and drug entities, including drug mechanisms and disease descriptions, we additionally integrate the textual details from multiple databases for gene/protein and pathway entities such as genomic position, gene activity summary and pathway orthologous event. We include the public sources for SKBs construction in Table 10 (Appendix B).
footnotetext: 1 We filter out non-English language papers as we only considers single-lingual queries.
@@ 2.3 Retrieval Tasks on Semi-structured Knowledge Bases
STaRK-Prime 109 43 2.61
We develop three novel retrieval-based question-answering datasets. To highlight, the queries feature the synthesization between relational and textual knowledge. Moreover, to elevate their applicability in practical scenarios, these queries are designed to mimic real-life query patterns in terms of the natural-sounding property and flexible formats. We detail our retrieval benchmarks’ scale in Table 3 and introduce the specifics of each retrieval dataset as follows:
STaRK-Amazon. This dataset comprises synthesized 9,100 customer-focused queries along with 81 human-generated queries for additional evaluation. The task aims at product searches, with a notable 68% of the synthesized queries yielding multiple answers. The dataset prioritizes customer-oriented criteria, highlighting textual elements such as product quality, functionality, and style. Additionally, it accentuates relational aspects, including brand, category, color and product connections (e.g., complementary or substitute items). The queries are framed in detailed, conversation-like formats, enriching the context and enhancing the dataset’s relevance to real-world scenarios.
STaRK-MAG. This dataset comprises 13,323 synthesized and 84 human-generated queries targeting at academic paper search. Beyond the single-hop relational requirements on STaRK-Amazon, STaRK-MAG also emphasizes the fusion between the textual requirements with multi-hop queries. For example, “Are there any papers from King’s College London” highlights the metapath (institution $\rightarrow$ author $\rightarrow$ paper) on the relational structure. We designed three single-hop and four multi-hop relational query templates, cf. Appendix B.2. The textual aspects of the queries diversifies, focusing on different elements such as the paper’s topic, methodology, or the advancements it introduces, sourced mainly from the abstracts.
STaRK-Prime. This dataset, comprising 11,204 synthesized queries and 109 human-generated queries across all ten entity types, incorporates single-hop and multi-hop relational information similar to STaRK-MAG. For synthesized queries, we developed 28 multi-hop query templates, detailed in Appendix B.3, to cover various relation types and ensure their practical relevance. For instance, the query “What is the drug that targets the genes or proteins expressed in <anatomy>?” serves applications in precision medicine and pharmacogenomics, aiding researchers and healthcare professionals in identifying drugs that act on genes or proteins associated with specific anatomical areas and enabling more targeted treatments. For entities like drug, disease, gene/protein, and pathway, the queries are a hybrid of relational requirements and textual requirements drawn from documents. A notable feature is the deliberate simulation of three distinct roles – medical scientist, doctor, and patient – for certain queries related to drugs and diseases. This is intended to diversify the language used across various user types and to evaluate the robustness of the retrieval system under varied linguistic scenarios. For entities such as effect/phenotype, the queries rely solely on relational data due to limited textual information in the knowledge base.
We divide the synthesized queries on each dataset into training, validation, and testing subsets, with the ratios detailed in Table 3, while using the human-generated queries as additional sets for evaluation. For the STaRK-Amazon dataset, we randomly allocate a subset of queries with 20 or fewer answers to the validation and testing sets, while the remaining queries are assigned to the training set. For the STaRK-MAG and STaRK-Prime, we use random splits to divide the data.
@@ 2.4 Benchmark Construction: Synthesized Queries
Figure 3: The process of constructing semi-structured retrieval datasets involves four main steps: 1) Sample Relational Requirement: Based on relational templates, sample a relational requirement on a SKB. 2) Extract Textual Properties: From a node that meets the relational requirement, extract relevant textual properties. 3) Combine Information: Merge the relational information and textual properties to form a natural-sounding query. 4) Construct Ground Truth Nodes: Check if nodes satisfy the textual properties using multiple language models to establish ground truth nodes.
@@
We present the novel pipeline that synthesizes the large-scale retrieval datasets. The core idea is to entangle relational information and textual properties and accurately construct ground truth answers. The construction of the retrieval datasets (Figure 3) generally involves four steps, and the specific processes vary depending on the characteristics of each dataset. These steps are as follows:
  • 1) Sample Relational Requirements: Initially, we sample a relation template, such as “(a product) belongs to <brand>” and ground it into a relational requirement, e.g., “belongs to Radio Flyer” in Figure 3. Each relational requirement yields a set of candidate entities that meet the requirement, i.e., products belonging to Radio Flyer. The relation templates for the three SKBs are constructed with expert/domain knowledge, which are available in Appendix B. Each relation template in the SKBs serves a distinct and practical purpose. For instance, the template “What is the drug that targets the genes or proteins which are expressed in <anatomy>?” is particularly valuable for precision medicine inquiries. It aids in pinpointing medications aimed at specific genetic or protein targets within defined anatomical areas, thereby contributing to the development of targeted treatments.
@@
  • 2) Extracting Textual Properties: The relational requirement yields a set of satisfying candidates. Subsequently, we select one of the candidate nodes, referred to as the gold answer, and use LLMs to extract textual properties of interest from the its document. We extract the textual properties that align with the interests of specific roles, such as customers, researchers, or medical scientists, according to the specific SKB. For example, in Figure 3, we extract the phrases of functionality and user experience from a Radio Flyer product from different locations of the product document.
  • 3) Combining Textual and Relational Information: After obtaining the textual and relational requirements, we synthesize them into a query using LLMs. We conduct two-stage synthesization using two different LLMs to avoid bias that might arise from relying on a single LLM and ensure a more diverse set of simulated queries. Specifically, this first-stage synthesization is guided by various criteria, such as natural sounding, adhering to the style of ArXiv searches, or aligning with specific roles. Furthermore, the second stage instructs the LLM to enrich the context and rephrase the language, demanding reasoning abilities in comprehending the requirements of the query.
  • 4) Filtering Additional Answers: Finally, we employ multiple LLMs to verify if each remaining candidate, excluding the gold answer used for extracting textual properties, meets the additional textual requirement. Only candidates that pass the verification across all LLMs are included in the final ground truth answer set. This stringent verification ensures the accuracy of the nodes in the ground truth answers. To validate the precision of this filtering process, we compute the average precision of each gold answer passes the verification. The precision rates on STaRK-Amazon, STaRK-MAG, and STaRK-Prime are 86.6%, 98.9%, and 92.3%, respectively, demonstrating the effectiveness of our filtering approach in maintaining high-quality ground truth answers.
@@
This dataset construction pipeline is automatic, efficient, and broadly applicable to the SKBs in our formulation. We include all of the prompts and the LLMs versions in the above steps in Appendix C.
2.5 Synthesized Data Distribution Analysis and Human Evaluation
@@
Table 4: Query diversity measurement on STaRK. See Appendix A for the metric definition.
Shannon Type-Token
Entropy Ratio
@@
We study the data distribution to understand the characteristics of our benchmark datasets and conduct human evaluation on the benchmark quality. We study the data distribution in three dimensions:
  • Query and Answer Length. We analyze the query length distribution (measured in the number of words) and the number of ground truth answers for each query. The query length reflects the amount of context information provided by users, while the number of ground truth answers indicates the level of inclusiveness or ambiguity of queries within the specific SKB. As illustrated in Figure 4, all three datasets exhibit similar query length distributions, with most queries containing around 16 words. Queries with up to 50 words typically involve mentions of other entities, such as product or paper titles, or provide more detailed context, such as specific symptoms descriptions. Interestingly, the answer length distribution for STaRK-Amazon shows a more significant long-tail pattern, with approximately 22% of answers exceeding 30 entities, whereas the answer lengths for STaRK-Prime and STaRK-MAG are all within 20 entities. On average, as shown in Table 3, STaRK-Amazon presents an average answer length of 5.32, indicative of the e-commerce recommendation domain where general customer inquiries frequently result in diverse recommendations. While STaRK-Prime has the smallest average answer length, partially attributed to the smaller size of the Prime SKB compared to the other two SKBs.

### v3

@@ Abstract
Answering real-world complex queries, such as complex product search, often requires accurate retrieval from semi-structured knowledge bases that involve blend of unstructured (e.g., textual descriptions of products) and structured (e.g., entity relations of products) information. However, many previous works studied textual and relational retrieval tasks as separate topics. To address the gap, we develop STaRK, a large-scale Semi-structure retrieval benchmark on Textual and Relational Knowledge Bases. Our benchmark covers three domains: product search, academic paper search, and queries in precision medicine. We design a novel pipeline to synthesize realistic user queries that integrate diverse relational information and complex textual properties, together with their ground-truth answers (items). We conduct rigorous human evaluation to validate the quality of our synthesized queries. We further enhance the benchmark with high-quality human-generated queries to provide an authentic reference. STaRK serves as a comprehensive testbed for evaluating the performance of retrieval systems driven by large language models (LLMs). Our experiments suggest that STaRK presents significant challenges to the current retrieval and LLM systems, highlighting the need for more capable semi-structured retrieval systems.
@@
Table 1: STaRK QA examples which involve semi-structured (relational and textual) information.
@@ 1 Introduction
To answer such queries, the underlying knowledge can be represented in semi-structured knowledge bases (SKBs) [36, 51, 41], which integrate unstructured data, such as natural language descriptions and expressions (e.g., description of the tricycle), with structured data, like entity interactions on knowledge graphs (e.g., a tricycle “brand” is Radio Flyer). This allows the SKBs to represent comprehensive knowledge in specific applications, making them indispensable in domains such as e-commerce [16], social media [32], and precision medicine [9, 19, 24].
@@
Limitations of prior works. Prior works focused on either purely textual queries on unstructured knowledge [21, 15, 26, 30, 54, 13, 25, 56] or structured SQL [61, 60, 61, 60] or knowledge graph queries [5, 58, 46, 47, 59, 48, 8, 14, 17, 3], which are limited in the span of knowledge and inadequate to study the complexities of retrieval on SKBs. Recently, large language models (LLMs) have demonstrated significant potential on information retrieval tasks [62, 15, 31, 44]. Nevertheless, it remains an open question of how effectively LLMs can be applied to the challenging retrieval tasks on SKBs. Moreover, the existing works mainly focus mainly on general knowledge, e.g., from Wikipedia. However, the knowledge may commonly come from private sources, requiring retrieval systems to operate on private SKBs. Therefore, there is a gap of how current LLM retrieval systems handle the complex textual and relational requirements in queries that can involve private knowledge.
@@
Present work. To address this gap, we present a large-scale Semi-structure retrieval benchmark on Textual and Relational Knowledge Bases (STaRK) (Figure 1). The key technical challenge that we solve is how to accurately simulate user queries on SKBs. This difficulty arises from the interdependence of textual and relational information, which leads to challenges in precisely construct the ground-truth answers from millions of candidates. Additionally, ensuring that queries are useful and resembles real-world scenarios adds further complexity to the benchmarking process.
@@
We develop a novel pipeline that simulates user queries and constructs precise ground truth answers using three SKBs built from extensive texts and millions of entity relations from public sources. We validate the quality of queries in our benchmark through detailed analysis and human evaluation, focusing on their naturalness, diversity, and practicality. Furthermore, we incorporate 274 human-generated queries to compare with synthesized queries and enrich the testing scenarios. With STaRK, we delve deeper into retrieval tasks on SKBs, evaluate the capability of current retrieval systems, and provide insights for future advancement. Key features of STaRK are:
@@
  • Natural-sounding queries on SKBs (Table 1): Queries in our benchmark incorporate rich relational information and complex textual properties. Additionally, these queries closely mirror the types of questions users would naturally ask in real-life scenarios, e.g., with flexible query formats and possibly with additional contexts.
@@
  • Diverse domains: Our benchmark spans three knowledge bases***Explore the SKBs at https://stark.stanford.edu/skb_explorer.html for applications including product recommendation, academic paper search, and precision medicine inquiries. STaRK provides a comprehensive evaluation of retrieval systems across diverse contexts and domains.
We conduct extensive experiments on LLM retrieval systems, highlighting challenges in handling textual and relational data and latency on large-scale SKBs with millions of entities or relations. Finally, we offer insights into building more capable retrieval systems to handle real-world complexity.
@@ 2 Benchmarking Retrieval Tasks over Textual and Relational Knowledge
2.1 Problem Definition
We are given a Semi-Structured Knowledge Base (SKB), which consists of a knowledge graph $G$ and a collection of free text documents $D$. Formally, let $G=(V,E)$ be the knowledge graph, where $V$ is the set of nodes and $E\subseteq V\times V$ is the set of edges representing relationships between nodes. $D=\bigcup_{i\in V}D_{i}$ be the collection of free-form text documents associated with the nodes, where $D_{i}$ is the set of documents associated with node $i$. For example, the product knowledge graph in e-commerce can capture relationships between products and brands/colors/categories, and the corresponding text documents include product descriptions, reviews, etc.
2.2 Semi-structured Knowledge Bases (SKBs)
@@
As shown Figure 2, we construct three large-scale SKBs with the relational and textual information with each entity. See Table 2 for the basic data statistics and Appendix A.1 for details.
@@
Amazon Semi-structured Knowledge Base. The SKB features four entity types: product, brand, color, and category, and five relation types: also_bought, also_viewed between product entities, and has_brand/color/category associated with the products. We derive the textual information of product nodes by combining Amazon Product Reviews [16] with Amazon Q&A Data [33]. This provides a rich amount of texts, including product descriptions and customer reviews. For other entities, we extract their names or titles as the textual attributes. Amazon SKB features an extensive textual data largely contributed from customer reviews and Q&A.
@@
MAG Semi-structured Knowledge Base. This SKB includes node entities of paper, author, institute, and field_of_study. We derive its relational structure by extracting a subgraph from obgn-mag [20], which contains shared paper nodes with obgn-papers100M [20] and all non-paper nodes. We filter out non-English language papers as we only consider single-lingual queries. The paper documents include their titles and abstracts. Additionally, we integrating details from the Microsoft Academic Graph database (version 2019-03-22) [51, 45], providing extra textual information like paper venue, author and institution names. This SKB demonstrates a large number of relations associate with paper nodes, especially on citation and authorship relations.
@@
Prime Semi-structured Knowledge Base. We leverage the exisiting knowledge graph PrimeKG [9] which contains ten entity types including disease, gene/protein, and eighteen relation types, such as associated_with, indication. Compared to the Amazon and MAG SKBs, Prime SKB is denser and features a greater variety of relation types. While PrimeKG provides text information on disease and drug entities, we integrate the data from multiple databases for gene/protein and pathway entities such as genomic position, gene activity summary and pathway orthologous event.
@@ 2.3 Retrieval Tasks on Semi-structured Knowledge Bases
Our retrieval benchmark (Table 3) consists of three novel retrieval-based question-answering datasets, each comprising synthesized train/val/test sets with $9$k to $14$k queries in total and a high-quality human-generate query set. The queries synthesize relational and textual knowledge, mirroring real-world queries in terms of natural-sounding property and flexible formats.
STaRK-Amazon. The task aims at product recommendation, with a notable 68% of the synthesized queries yielding more than one ground truth answer. The dataset prioritizes customer-oriented criteria, highlighting textual elements such as product quality, functionality, and style. Moreover, it incorporate single-hop relational aspects (Appendix A.2) into the queries, including brand, category, and product connections (e.g., complementary or substitute items). The queries are framed in conversation-like formats, enriching the context and enhancing the dataset’s relevance to real-world scenarios.
STaRK-MAG. Beyond the single-hop relational requirements in STaRK-Amazon, STaRK-MAG emphasizes the fusion between the textual requirements with multi-hop queries for precise academic paper search. For example, “Are there any papers from King’s College London” highlights the metapath (institution $\rightarrow$ author $\rightarrow$ paper) on the relational structure. We designed three single-hop and four multi-hop relational query templates (Appendix A.3). The textual aspects focus on the paper’s topic, methodology, and contribution etc.
STaRK-Prime. The task is to answer complex biomedicine inquiries. For synthesized queries, we developed 28 multi-hop query templates (Appendix A.4) to cover various relation types and ensure their practical relevance. For example, the template “What is the drug that targets genes or proteins in <anatomy>?” aids precision medicine by identifying treatments targeted to specific anatomical areas. For drug, disease, gene/protein, and pathway entities, the queries are a hybrid of relational and textual requirements. For entities such as effect/phenotype, the queries rely solely on relational data due to limited textual information. We exhibit three distinct user roles – medical scientist, doctor, and patient – for generating queries about drug and disease, which diversify the language to comprehensively evaluate the retrieval systems.
@@
STaRK-Prime 98 41 2.77
Figure 3: The construct pipeline to generate our semi-structured retrieval datasets.
@@ 2.4 Benchmark Construction: Synthesized Queries
In Figure 3, we present a novel pipeline that synthesizes the SKB queries and automatically generates the ground truth answers. The key idea is to entangle relational and textual information during synthesis and disentangle them during answer filtering. It involves four steps as follows:
  • 1) Sample Relational Requirements: For each query, we sample a practical relation template constructed with expert/domain knowledge, e.g., (a product) belongs to <brand>” and ground it with sampled entities (i.e., a specific brand), e.g., “belongs to Radio Flyer”. This relational requirement yields a set of candidate entities, i.e., products belonging to Radio Flyer.
@@
  • 2) Extracting Textual Properties: We randomly sample a candidate entity from the first step, referred to as the gold answer, from which LLMs extract properties that align with the interests of specific roles (e.g., customers, researchers, or doctors) in its textual document. In Figure 3, we extract multiple properties about the functionality and user experience from a Radio Flyer product.
  • 3) Combining Textual and Relational Information: We use two LLMs to synthesize queries from textual properties and relational requirements, enhancing diversity and reducing bias arise from relying on a single LLM. The first LLM focuses on generating natural, role-specific, and style-consistent (e.g., ArXiv searches) queries. The second LLM enriches the context and rephrases queries, which poses the need for advanced reasoning to comprehend them under complex contexts.
  • 4) Filtering Additional Answers: Finally, we employ multiple LLMs to verify if the candidates from the first step meet the extracted textual properties. Only candidates passing all LLM verifications are included in the final ground truth set. To assess the precision of this filtering mechanism, we compute the average ratios for the gold answers to be verified, which are 86.6%, 98.9%, and 92.3% on the three datasets, highlighting our efficacy in yielding high-quality ground truth answers.
@@
This dataset construction pipeline is automatic, efficient, and broadly applicable to the SKBs in our formulation. We include all of the prompts and the LLMs versions in the above steps in Appendix E.
@@
Table 4: Query diversity measurement on STaRK. See Appendix B for the metric definition.
Shannon Entropy Type-Token Ratio
@@
2.5 Synthesized Data Distribution Analysis and Human Evaluation
  • Query and Answer Length. Query length (in words) reflects the amount of user-provided context information, while the number of answers indicates query ambiguity/concreteness. Figure 4 shows similar query length distributions across the datasets, with most queries around 16 words. Longer queries (up to 50 words) often mention other entities or provide detailed context. Notably, STaRK-Amazon has a significant long-tail pattern, with about 22% of the answers have more than 30 entities, reflecting diverse e-commerce recommendations and ambiguous user queries.